Find the value of dy when x=3 and dx= 0.01 for f(x)= 5x2-3. Compare this with the value of Ay when x=3 and Ax 0.01
Answer #1

f(x) = 5x²-3 6 f(x)= lox dy dr = lox dy - lox. da 2 = 10 (3)* .01 - 0.3 f(x+ox) -f(x) oy= f(3.01) +(3) 5(3) 5(3.01) ² - 3 0.3